OpenStudy (lukecrayonz):

Differentiate implicitly to find d^2y/dx^2 x^3-3y^3=4

OpenStudy (callisto):

1) Differentiate both sides w.r.t. x. What do you get?

OpenStudy (lukecrayonz):

3x^2 and not sure what wrtx means :P

OpenStudy (lukecrayonz):

with relations to x?

OpenStudy (callisto):

w.r.t = with respect to. Actually, you need to differentiate BOTH sides. \[\frac{d}{dx}(x^3-3y^3)=\frac{d}{dx}(4)\]

OpenStudy (callisto):

Consider the left side: \[\frac{d}{dx}(x^3-3y^3)=\frac{d}{dx}(x^3) - \frac{d}{dx}3y^3 = ...?\]

OpenStudy (b87lar):

One point to keep in mind is that y is a function of x. So you need to differentiate it as such (as opposed to viewing it as a constant and dropping it from the derivative)

OpenStudy (lukecrayonz):

It's not just simply moving everything so y is by itself?

OpenStudy (lukecrayonz):

such as 4-x^3=-3y^3 dividing 3, getting the cube root then finding the derivative of that?

OpenStudy (callisto):

If you make y as the subject, I guess it is no longer "differentiate implicitly"

OpenStudy (b87lar):

that would be an explicit differentiation. Sometimes functions cannot be expressed explicitly (thus they are also called implicit functions)

OpenStudy (b87lar):

so in an implicit differentiation you apply the chain rule to f(y(x)), such as in this case: (3y^3)'=(9y^2)y'

OpenStudy (lukecrayonz):

Makes sense, so d^2y/dx^2 is just implicit differentiation? Not double implicit?

OpenStudy (b87lar):

you will need to differentiate twice: both steps are implicit differentiation

OpenStudy (b87lar):

then you identify the term d^2y/dx^2 and solve for it (which is what they ask for)

OpenStudy (lukecrayonz):

So what is the final answer? Taking my final tomorrow and teacher never really went over this, want to quickly brush up on it

OpenStudy (lukecrayonz):

2x/3y^2?

OpenStudy (b87lar):

This is what i am getting: \[d^2/dx^2(x^3-3 y(x)^3 = 4\]\[6 x - 9 y(x) (y(x) y''(x)+2 y'(x)^2)=0\] Now you have to reshuffle things to solve for y''(x)

OpenStudy (b87lar):

The first derivative is\[y'(x) = x^2/(3 y(x)^2)\] (note I am using y(x) to emphasize y is a function of x, i hope this is not confusing)

OpenStudy (kc_kennylau):

For example, here is how you would differentiate x^2+y^2=1 implicitly: \[x^2+y^2=1\]\[\frac d{dx}(x+y)=\frac d{dx}1\]\[2x+\frac d{dx}y^2=0\]\[2x+\frac d{dy}y^2\frac{dy}{dx}=0\hspace{20pt}\mbox{(Chain rule)}\]\[2x+2y\frac{dy}{dx}=0\]\[\frac{dy}{dx}=-\frac yx\]

OpenStudy (kc_kennylau):

When you meet a function of y, you used the chain rule: 1. Differentiate the function wrt y 2. Multiply the derivative by \(\dfrac{dy}{dx}\)

OpenStudy (anonymous):

@kc_kennylau The last step of your example should read \(\large\dfrac{dy}{dx} = -\dfrac{x}{y}\). ;)

OpenStudy (kc_kennylau):

oh yes sorry

OpenStudy (lukecrayonz):

Question about this problem! http://gyazo.com/e2d41376f33f06aae1103cd8748fc7a4 Essentially it's just f'(x)=(3x^4+6)^3 when x=-1?

OpenStudy (kc_kennylau):

It's not f'(x) that is (3x^4+6)^3, it's f(x). (However what we're finding is f'(x))

OpenStudy (lukecrayonz):

I know :O But after substituting in 3x^4+6 for u, it's what I said?

OpenStudy (kc_kennylau):

yep